An electron is accelerated to a speed of 4.23 × 10^6 ms –1 from rest by a potential difference. Calculate the potential difference


1
Expert's answer
2021-12-22T14:11:51-0500

"e = 1.6*10^{-19}C"

"v = 4.23*10^6 \frac{m}{s}"

"m = 9.1*10^{-31}kg"

"\Delta \phi =" ?


"A=\Delta E"

"e\Delta \phi = \large\frac{mv^2}{2} \to \Delta \phi = \large\frac{mv^2}{2e}=\frac{9.1*10^{-31}*(4.23)^2*10^{12}}{2*1.6*10^{-19}}=51V"